OpenStudy (anonymous):

can someone help me please: solve x=1-2y 3x-y=17 by substitution techniche

OpenStudy (anonymous):

substitute the first equation in the second one. 3(1-2y)-y=17 3-6y-y=17 -7y=14 y=-2 substitute back y in the first equation x=1-2(-2) x=1+4 x=5

OpenStudy (anonymous):

how did you get -7

OpenStudy (anonymous):

girl : Am I pretty? Boy : NO. Girl : Do you want to be with me forever? Boy : NO. Girl : Would you cry if I walked away?Boy : NO. She heard enough, and was hurt. She walked away, tears ran down her face. The boy grabbed her arm. Boy : Your not pretty, your beautiful. I don't want to be with you forever, I NEED to be with you forever. And I wouldn't cry if you walked away, I would DIE. (Boy whispers) : Please? Stay with me. (Girl whispers) : I will. Tonight at midnight your true love will realize they love you. Something good will happen to you between 1-4 pm. Tomorrow it could be anywhere. Get ready for the biggest shock of your life! If you don't post this to 5 other pages. You will have relationship problems for the next 10 years